# Pulse-train propagation in nonlinear Kerr media governed by higher-order dispersion

### v1

@@ 1. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{dn}\left(x,k\right)$-type
We introduce an appropriate ansatz for the propagating waveform solution of Eq. (7) as,
$\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{dn}\left(w(\xi-\xi_{0}),k\right),$ (9)
where $\xi_{0}$ is an arbitrary constant while $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)$ are the cosine and third kind of Jacobian elliptic functions of modulus $k$ taking values $10<k<$ 1.
Substituting the ansatz (9) into Eq. (7) and setting the coefficients of independent terms equal to zero, we obtain the following algebraic equations:
@@
$w=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{5\epsilon^{2}(k^{2}+1)}},~{}~{}~{}~{}P=\pm\frac{k}{k^{2}+1}\sqrt{\frac{-6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (13)
@@
$c=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right)^{2},$ (14)
The wave number $\kappa$ follows from Eqs. (8) and (14) as
@@
Incorporating these results into Eq. (2), we obtain the following periodic wave solution to the extended NLSE (1):
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{dn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (16)
where $\zeta=\tau-v^{-1}z-\xi_{0}$, with $\xi_{0}$ being the position of the wave at $z=0$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
Physically, Eq. (16) describes a periodic wave with amplitude $P$ and inverse temporal width $w$ depending on all order of dispersion as well as nonlinearity.
@@
$P\tau_{0}^{2}=-2k\left(-\Omega\right)^{1/2}.$ (17)
where the parameter $\Omega=30\epsilon/\gamma$ has been introduced for brevity.
@@ 2. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
where $\mathit{\mathrm{sn}}$ is Jacobi elliptic sine function of modulus $k$ which is in the range $10<k<$ 1.
Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
@@
$0120\epsilon\eta^{4}k^{2}-\gamma P^{2}=$ 0 . (21)
Thus we have the following relations for the pulse inverse width $w$ and amplitude $\lambda$,
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(4-3k^{2})}},~{}~{}~{}~{}P=\pm\frac{k}{4-3k^{2}}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (22)
with $\gamma\epsilon>0$ and $3\sigma^{2}>8\alpha\epsilon$. Additionally, Eqs. (18) and (21) yield the parameter $c$as
$c=\frac{b^{2}(244k^{2}-89k^{4}-144)}{100\epsilon(4-3k^{2})^{2}},$ (23)
Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
@@
These equations yield the periodic wave solution as
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (25)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Noting that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by Eqs. (5) and (6), respectively.
@@ 3. Pulse-trains of the $\mathrm{dn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
Next we make an ansatz for Eq. (7) as,
$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (27)
Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
@@
$0120\epsilon w^{4}k^{2}-\gamma P^{2}=$ 0 . (30)
Thus we have the following relations for the pulse inverse width $\mu$ and amplitude $R$,
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(2k^{2}-1)}},~{}~{}~{}~{}P=\pm\frac{k}{2k^{2}-1}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (31)
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ when $11/\sqrt{2}<k<$ 1 and for $8\alpha\epsilon>3\sigma^{2}$ when $0<k<1/\sqrt{2}$.
Moreover, Eqs. (18) and (21) yield the parameter $c$as
$c=\frac{b^{2}(64k^{2}-64k^{4}+11)}{100\epsilon(2k^{2}-1)^{2}},$ (32)
Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
@@
These equations yield the periodic wave solution as
$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (34)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
@@
On comparing Eqs. (16), (25) and (34), we infer that under the influence of second-, third-, and fourth-order dispersions and Kerr nonlinearity the fiber medium allows the propagation of distinct types of pulse-trains which present a relation between the pulse amplitude and duration determined by the sign of the joint parameter $\Omega=30\epsilon/\gamma$ only. To our knowledge, the pulse-trains obtained above for the extended NLSE (1) are firstly reported in this paper.
@@
Figure 1(a) depicts the intensity profile of the pulse-train solutions (16) of Eq. (1) for the parameters values: $\alpha=-1,$ $\gamma=2,$ $\sigma=-0.25,$ and $\epsilon=-0.25.$ The results for the pulse-train solutions (25) and (34) are illustrated in Figs. 1(b) and 1(c) for the values $\alpha=0.49,$ $\gamma=2,$ $\sigma=1,$ and $\epsilon=0.5$. Here the position $\xi_{0}$ of the waveform solutions at $z=0$ is choosen to be equal to zero. Also, the value of the elliptic modulus $k$ is taken as $k=0.9$. In Fig. 2, we show the evolution of these pulse-train solutions in the fiber medium. From these figures, we can see that all these pulse-trains are positive and exhibit an oscillating character. In addition, we can see that the pulse-train solutions (25) and (34)  present two peaks in each periodic unit, markedly different from the pulse train solution (16) which has a one peak structure.
@@
It is also interesting to see from Figs. (3) and (4) that the shape of the pulse-train solution (25) is similar to that for the pulse-train solution (34) for the value of the elliptic modulus $k=0.999$.
@@
The key observation is that the amplitudes, widths and wavenumbers of the derived pulse-train solutions are different whereas their velocity is the same. More importantly, the relation (6) shows that the velocity of the pulse-trains can be significantly reduced for suitable parameters of the waveguiding media, which may find application in developing slow light systems. In Fig. 5, we present the dependence of velocity on the group velocity dispersion parameter $\alpha$ of the fiber for three different values of the quartic dispersion parameter $\epsilon$. It can clearly be seen that the velocity decreases intensely in the magnitude with the increase of dispersion parameter $\alpha$. The decrease of velocity implies that the slowing motion of pulse-trains is possible in the transmitting medium when considering large dispersion parameter $\alpha$. It is also observed that the pulse velocity has large magnitude for small negative values of $\epsilon$. Hence, one concludes that the velocity of pulse-trains can be controlled by judicious choice of the dispersion parameters.
@@ III.2 Quartic and dipole soliton solutions
$\psi(z,\tau)=P_{0}~{}\mathrm{sech}^{2}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (36)
@@
$\kappa=-\frac{4}{25\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (39)
with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Also the velocity $v$ and frequency shift $\delta$ in this quartic soliton solution are defined by Eqs. (5) and (6).
In addion, when $k\rightarrow 1$, the periodic solution (25) transforms into our dipole soliton solution [7]:
$\psi(z,\tau)=P_{0}~{}\mathrm{sech}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\mathrm{th}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (40)
where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse take the form
@@
$\kappa=\frac{11}{100\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (43)
@@
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$. The velocity $v$ and frequency shift $\delta$ in this dipole soliton solution are given by Eqs. (5) and (6).
It is interesting to note that if one considers the limiting case with $k=1$, the periodic wave solution (34) also degenerates to the same soliton-like solutions (40).
IV Conclusion

### v2

@@ 1. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{dn}\left(x,k\right)$-type
We introduce an appropriate ansatz for the propagating waveform solution of Eq. (7) as
$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{cn}\left(w(\xi-\xi_{0}),k\right),$ (9)
with cosine and the third kind of Jacobian elliptic functions $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)$ of the modulus number $k$ ($0<k<1$). Here $P$ is the amplitude of the wave to be determined and $\xi_{0}$ is its initial position (at $z=0$).
Substituting the ansatz (9) into Eq. (7) and setting the coefficients of $\mathrm{dn}^{j}\left(w(\xi-\xi_{0}),k\right)\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ functions (where $j=1,3,5$) equal to zero, one gets the algebraic equations:
@@
$w=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{5\epsilon^{2}(k^{2}+1)}},~~~~P=\pm\frac{k}{k^{2}+1}\sqrt{\frac{-6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (13)
@@
$c=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right)^{2}.$ (14)
Then using Eqs. (8) and (14), we can obtain the expression for the wave number $\kappa$:
@@
Based on these findings, we get a novel periodic wave solution for the model (1) as
$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{cn}\left(w\zeta,k\right)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (16)
where $\zeta=\tau-v^{-1}z-\xi_{0}$. Note that the frequency shift $\delta$ and velocity $v$ of this periodic waveform are defined by the relations (5) and (6), respectively.
From Eq. (13), one can see that the amplitude $P$ and inverse temporal width $w$ of this periodic wave crucially depend on various dispersion effects as well as nonlinearity. This result indicates that the contributions of dispersions and self-phase modulation processes are essential for this newly found pulse-train to be formed in the optical fiber medium.
@@
$P\tau_{0}^{2}=-2k\left(-\Omega\right)^{1/2},$ (17)
where the parameter $\Omega=30\epsilon/\gamma$ has been introduced for brevity. Note that here, we considered the ++ sign in all the expressions of the wave amplitude.
@@ 2. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
with cosine and sine Jacobian elliptic functions $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{sn}\left(w(\xi-\xi_{0}),k\right)$ of the modulus number $k$ ($0<k<1$).
Upon substituting of this ansatz into Eq. (7) and then setting the coefficients of $\mathrm{cn}^{j}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right)$ functions (where $j=1,3,5$) equal to zero, we obtain the algebraic equations:
@@
$120\epsilon\eta^{4}k^{2}-\gamma P^{2}=0.$ (21)
Based on these equations, one obtains the relations for the inverse width $w$ of the wave and its amplitude $P$ as
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(4-3k^{2})}},~~~~P=\pm\frac{k}{4-3k^{2}}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (22)
with $\gamma\epsilon>0$ and $3\sigma^{2}>8\alpha\epsilon$. Additionally, by using Eqs. (19) and (22), we have the expression:
$c=\frac{b^{2}(244k^{2}-89k^{4}-144)}{100\epsilon(4-3k^{2})^{2}}.$ (23)
Then from Eqs. (8) and (23), we find that the wave number $\kappa$ of the periodic wave solution takes the form as
@@
These results yield a new kind of pulse-train solution for the NLSE model (1) as
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (25)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Noting that the frequency shift $\delta$ and velocity $v$ in this periodic structure are defined by the relations (5) and (6), respectively.
@@ 3. Pulse-trains of the $\mathrm{dn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
Next we make an ansatz for Eq. (7) as
$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right).$ (27)
Substituting this ansatz into the amplitude equation (7) and then setting the coefficients of $\mathrm{dn}^{j}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right)$ functions (where $j=1,3,5$) equal to zero, one finds the algebraic equations:
@@
$120\epsilon w^{4}k^{2}-\gamma P^{2}=0.$ (30)
Solving these equations, one can get the pulse parameters $w$ and $P$ as
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(2k^{2}-1)}},~~~~P=\pm\frac{k}{2k^{2}-1}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (31)
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ when $1/\sqrt{2}<k<1$ and $8\alpha\epsilon>3\sigma^{2}$ when $0<k<1/\sqrt{2}$.
Moreover, the use of Eqs. (28) and (31) lead to an expression for the parameter cc\ as
$c=\frac{b^{2}(64k^{2}-64k^{4}+11)}{100\epsilon(2k^{2}-1)^{2}}.$ (32)
Now using Eqs. (8) and (32), we can determine the wave number $\kappa$ as
@@
Thus, we find a novel pulse-train solution for the NLSE model (1) as
$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (34)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Here $\delta$ and $v$ can be worked out by using the relations (5) and (6), respectively.
@@
On comparing Eqs. (17), (26) and (35), we infer that under the effects of quadratic, cubic and quartic dispersions and self-phase modulation nonlinearity, the fiber system allows the propagation of distinct types of pulse-trains which present a relation between the pulse amplitude and duration determined by the sign of the joint parameter $\Omega=30\epsilon/\gamma$ only. To our knowledge, the pulse-trains derived above for Eq. (1) are firstly presented in this work.
@@
Figure 1(a) depicts the intensity profile of the pulse-train solution (16) of the model (1) for the parameters values: $\alpha=-1,$ $\gamma=2,$ $\sigma=-0.25,$ and $\epsilon=-0.25.$ The results for the pulse-train solutions (25) and (34) are shown in Figs. 1(b) and 1(c) for the values $\alpha=0.49,$ $\gamma=2,$ $\sigma=1,$ and $\epsilon=0.5$. One notes here that the value of initial position $\xi_{0}$ (at $z=0$) of the nonlinear waves is chosen zero. Also, the modulus number $k$ is selected to be $k=0.9$. The evolution of these pulse-train solutions in the fiber medium are displayed in Fig. 2. From these figures, we observe that the three types of pulse-train solutions are positive and exhibit an oscillating behavior. In addition, we can see that the pulse-train solutions (25) and (34) present two peaks in each periodic unit, markedly different from the pulse train solution (16) which has a one peak structure.
@@
It is also interesting to see from Figs. (3) and (4) that the shape of the pulse-train solution (25) is similar to that for the pulse-train solution (34) for the modulus number $k=0.999$.
A key observation is that the amplitudes, widths and wave numbers of the derived pulse-train solutions are different whereas their velocity is the same. More importantly, the expression (6) demonstrates that the velocity of these pulse-trains can be considerably diminished for suitable parameters of the waveguiding media, which have application value in the development of slow-light systems. Recently, Li and Huang [48] studied the modification of a slow-light soliton in three-level atomic systems. They derived a high-order NLSE incorporating third-order dispersion, delay in nonlinear refractive index, modulation induced by the microwave field, nonlinear dispersion, and correction terms of linear and differential absorptions; and demonstrated that an obvious decrease of propagating velocity of the soliton can be obtained in the presence of the microwave field. Note that photonic crystal waveguides (PhCWs) represent interesting slow light structures due to their confinement and versatile dispersion properties including slow light and possible control of group velocity dispersion effects [49, 50, 51]. In order to strictly answer the question of possibility of applying the newly found pulse-trains in the development of slow-light systems much further research is needed. A detailed investigation of this problem will be reported in a forthcoming publication.
@@
In Fig. 5, we present the dependence of velocity on the group velocity dispersion parameter $\alpha$ of the medium for three distinct values of quartic dispersion coefficient $\epsilon$. We clearly see that the wave velocity decreases intensely in the magnitude as the value of dispersion parameter $\alpha$ is increased. The decrease of velocity implies that the slowing motion of pulse-trains is possible in the transmitting medium when considering large dispersion parameter $\alpha$. It is also observed that the pulse velocity has large magnitude for small negative values of $\epsilon$. Hence, one concludes that the velocity of pulse-trains can be controlled by judicious choice of the dispersion coefficients.
@@ III.2 Quartic and dipole soliton solutions
$\psi(z,\tau)=P_{0}~\mathrm{sech}^{2}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (36)
@@
$\kappa=-\frac{4}{25\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right),$ (39)
with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Here the soliton frequency shift $\delta$ is given by Eq. (5) and the velocity $v$ is defined by Eq. (6).
Notice that for this quartic soliton waveform, the relation between the amplitude and duration is given by
$P_{0}\tau_{0}^{2}=-2\sqrt{\frac{-30\epsilon}{\gamma}}.$ (40)
provided that $\gamma\epsilon<0$.
In addition, when $k\rightarrow 1$, the periodic solution (25) transforms into our dipole soliton solution [36]:
$\psi(z,\tau)=P_{0}~\mathrm{sech}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\mathrm{th}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (41)
where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse take the form,
@@
$\kappa=\frac{11}{100\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right),$ (44)
@@
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$. Notice that the frequency shift $\delta$ of this dipole soliton is given by Eq. (5) and its velocity $v$ is defined by Eq. (6).
It is interesting to note that if one considers the limiting case with $k=1$, the pulse-train solution (34) also becomes the same dipole-type soliton solution (41). Notice that for this kind of solitons, we obtain a relation between the pulse amplitude and duration as:
$P_{0}\tau_{0}^{2}=2\sqrt{\frac{30\epsilon}{\gamma}}.$ (45)
provided that $\gamma\epsilon>0$.
IV Propagation of pulse-trains in extended NLSE with higher-order dispersion
For completeness, we have numerically simulated the propagation of pulse-train solutions (16), (25) and (34) in Eq. (1) by utilizing the split-step Fourier method [52]. The physical parameters used in this model for the case of pulse-train solution (16) are representative of a femtosecond Ti: sapphire laser [53]: $\beta_{2}=-100$ fs2/cm, $\beta_{3}=-500$ fs3/cm, $\beta_{4}=-5000$ fs4/cm, and $\gamma=2.36\times 10^{-6}$ W-1cm-1. Note that these experimental parameters satisfy the relations $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$ which are necessary and sufficient for the existence of the pulse-train solution (16). As concerns the pulse-train solutions (25) and (34), we use the parameters of the silicon-based slot waveguide as [36, 41]: $\beta_{2}=-0.05$ ps2/m, $\beta_{3}=1.9\times 10^{-5}$ ps3/m, $\beta_{4}=2.5\times 10^{-5}$ ps4/m, and $\gamma=42$ W-1cm-1. We point out that with these system parameters, the relations $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ for the existence of pulse-train solutions (25) and (34) are fulfilled. Figure 6 illustrates the evolutions of periodic pulse-trains propagating through a distance of $20$ dispersion lengths. Here the elliptic modulus is chosen as $k=0.9$. We can see that the pulse-train solutions (16), (25) and (34) can preserve their shape during propagation due to the balance between the effects of quadratic, cubic and quartic dispersions and the Kerr nonlinearity in the system.
Finally, the figures presented above indicate that the newly found pulse-train solutions have no oscillated tails on the wings of pulses. It important to mention that experimental and numerical studies have demonstrated that solitons propagating in nonlinear Kerr media with dominant fourth-order dispersion are approximately Gaussian with oscillating tails [54, 55]. By considering the influence of both second-, third- and fourth-order dispersions on the pulse propagation through fibers, one of the present authors has recently obtained quartic solitons with a $\mathrm{sech}^{2}$ shape which does not carry oscillating tails [35]. Under these conditions, we have also found that dipole solitons without oscillating tails are formed in the highly dispersive fiber system [36]. These important results show that dispersive effects up to the fourth order and Kerr nonlinearity can achieve a perfect balance under certain conditions and lead to the formation of solitons without oscillating tails. No doubt, the pulse-trains presented here may be of guidance for the experimental generation of optical periodic waves due to their precise nature and interesting amplitude-duration relationship.
V Stability of the pulse-train solutions
[…]
Figure 7: The numerical evolution; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34) under the perturbation of white noise with maximal value as $0.1$. The values of parameters are the same as in Figs. 2 (a), (b) and (c) respectively.
VI Conclusion